# bibCyclic.bib

### v2

@@ Abstract
We study the $K$-theoretic enumerative geometry of cyclic Nakajima quiver varieties, with particular focus on $\text{Hilb}^{m}([\mathbb{C}^{2}/\mathbb{Z}_{l}])$, the equivariant Hilbert scheme of points on $\mathbb{C}^{2}$. The direct sum over $m$ of the equivariant $K$-theories of these varieties is known to be isomorphic to the ring symmetric functions in $l$ colors, with structure sheaves of torus fixed points identified with wreath Macdonald polynomials. Using properties of wreath Macdonald polynomials and the recent identification of the Maulik-Okounkov quantum affine algebra for cyclic quivers with the quantum toroidal algebras of type $A$, we derive an explicit formula for the generating function of capped vertex functions of $\text{Hilb}^{m}([\mathbb{C}^{2}/\mathbb{Z}_{l}])$ with descendants given by exterior powers of the 00th tautological bundle. We also sharpen the large framing vanishing results of Okounkov, providing a class of descendants and cyclic quiver varieties for which the capped vertex functions are purely classical.
@@ 1.1 Enumerative geometry of quiver varieties
The pioneering work of Maulik-Okounkov [MO] demonstrated that the quantum cohomology of Nakajima quiver varieties is controlled by certain geometrically constructed quantum groups. The $K$-theoretic version of this work, relating quasimaps counts to quantum affine algebras, was developed in [Ok17] and [OkSm]. This interplay between geometry and representation theory has lead to a large body of work in areas like symmetric functions, integrable systems, characteristic classes, and $q$-difference equations, [AOElliptic, AgOk, BRbows, NegGor, KorZeit2, indstab2, OkSm, MirSym2, MirSym1, qdehilb, zhu2]. On the one hand, $K$-theoretic counts present additional technical challenges compared to cohomological counts; on the other hand, they are required to observe additional structure, most importantly 3d mirror symmetry [AOElliptic, AgOk, dinkms2, dinkms1, dinksmir2, dinksmir]. By the degeneration and glue formulas [Ok17], general $K$-theoretic quasimap counts in Nakajima quiver varieties are built from the following three objects: the (bare) vertex, the capping operator, and the capped vertex. Each corresponds to a different choice of “boundary condition” for quasimaps. We touch briefly on the the first two before turning in more detail to the third, which is the main focus of this paper.
@@
For many examples of interest, the bare vertex can be computed by localization, see [AgOk, dinksmir2]. It provides a class of hypergeometric series generalizing the well-known ${}_{n+1}\phi_{n}$ basic hypergeometric series studied at length in, for example, [hypergeo]. Vertex functions are known to satisfy certain $q$-difference equations with regular singularities. The monodromy of these equations was determined in [AOElliptic] and [indstab2].
@@
The capping operator is the fundamental solution of the quantum difference equation, which is the $K$-theoretic version of the quantum differential equation. For Nakajima quiver varieties, these equations generalize the qKZ and dynamical difference equations studied in [FR, TV3, TV2]. The geometric approach to these equations was used by Aganagic-Okounkov in [AgOk] to provide integral formulas for solutions.
@@
The third fundamental curve count, the capped vertex, is the main object of interest for this paper. In some ways, the capped vertex is simpler than the first two curve counts: it is defined in non-localized $K$-theory and is purely classical for quiver varieties with large enough framing, a phenomenon known as “large framing vanishing” and proven in [Ok17]. It was proven by Smirnov [Sm16] for the Jordan quiver and by Zhu [zhu3] for general quivers (and earlier in cohomology in [PPRationality]) that the capped vertex is a rational functions of the quantum parameters (typically called Kähler parameters in this setting). On the other hand, computing the capped vertex explicitly in examples is difficult; the localization theorem essentially reduces the problem to computing both the bare vertex and capping operator. Building on [Sm16], the recent work of Smirnov and the first author [AySm] developed explicit formulas for the capped vertex for what is arguably the most important quiver variety: the Hilbert scheme of points in the plane. Their formulas reveal a beautiful relationship with Macdonald polynomials and give an enumerative-geometric interpretation to certain evaluations of them.
@@
In this paper, we push the methods of Ayers-Smirnov further, studying the capped vertex functions for cyclic quiver varieties. Our main theorem, Theorem 1.1 below, describes the capped vertex functions for certain cyclic quiver varieties in terms of wreath Macdonald polynomials [BFwreath, Haiman1]. Our main theorem can be interpreted as describing capped vertex functions as certain evaluations of wreath Macdonald polynomials, see Theorem 1.2.
@@ 1.2 The capped descendant vertex
Let $\mathcal{M}:=\text{Hilb}^{n}([\mathbb{C}^{2}/\Gamma_{l}])$ where $\Gamma_{l}=\mathbb{Z}/l\mathbb{Z}$, which acts on $\mathbb{C}^{2}$ by rotations. Concretely, the variety $\mathcal{M}$ parameterizes $\Gamma_{l}$ invariant ideals $I\subset\mathbb{C}[x,y]$ such that $\text{char}_{\Gamma_{l}}\left(\mathbb{C}[x,y]/I\right)=\sum_{i=0}^{l-1}n\chi^{i}$, where $\chi^{i}$ are the irreducible characters of $\Gamma_{l}$. In particular, $\mathcal{M}\subset\text{Hilb}^{nl}(\mathbb{C}^{2})$.
The variety $\mathcal{M}$ is an example of a Nakajima quiver variety [NakALE, NakQv] with dimension vector $\mathsf{v}=(n,n,\ldots,n)$ and thus has a canonical presentation as a geometric invariant theory quotient
@@
This description is reviewed in section 3. Hence we can apply the general machinery of quasimaps to GIT quotients developed in [qm]. To that end, let $\text{QM}^{d}$ be the moduli space of stable quasimaps of degree $d\in\mathbb{Z}^{l}$ relative to the point $0\in\mathbb{P}^{1}$. It compactifies the space of genuine maps from $\mathbb{P}^{1}$ to $\mathcal{M}$.
@@
By definition of $\mathcal{M}$, there is an inclusion $\iota:\mathcal{M}\hookrightarrow\mathcal{M}^{\text{st}}$ into the stack quotient. There is an action of a torus $\mathsf{T}$ on each piece of the above diagram such that the evaluation maps are $\mathsf{T}$-equivariant.
@@
The series $\hat{V}^{\left(\tau\right)}$ was proven in [Sm16] to be a rational function in $z$ for the rank $r$ instanton moduli space, for which the cyclic quiver at $l=1$ is the rank 1 version. An explicit formula for the $l=1$ case was given in [AySm]. Quasimap counts in cyclic quiver varieties are related to curve counts in threefolds, see [liu]. The cohomological version of this relationship goes back to Okounkov-Pandharipande [OPQuanCoh] and Maulik-Oblomkov [ObloMaul].
@@
where the superscript Sym refers to Laurent polynomials that symmetric in the Chern roots $x_{i,1},\ldots,x_{i,n}$ for each $i$ separately. Equivalently, $\mathcal{D}$ is the ring generated by the Schur functors of the tautological vector bundles $\mathcal{V}_{i}$, $10\leq i\leq l-$ 1, on $\mathcal{M}$. Let $\tau_{i}:=\sum_{j=0}^{n}\bigwedge^{j}\mathcal{V}_{i}$.
@@
Now we consider the varieties $\mathcal{M}$ for all $n$ and denote $\mathcal{M}(n):=\text{Hilb}^{n}([\mathbb{C}^{2}/\Gamma_{l}])$. By [BFwreath], there is an isomorphism
@@
where $\Lambda_{t_{1},t_{2}}=\mathbb{Q}(t_{1},t_{2})[p_{1},p_{2},\ldots]$ is the ring of symmetric functions [^2] The usual $q$ and $t$ of Macdonald theory match our parameters by $q=t_{1},t=t_{2}$.. The ring $\Lambda_{t_{1},t_{2}}^{\otimes l}$ is generated as an algebra over $\mathbb{Q}(t_{1},t_{2})$ by the “colored” power sums: $p^{(i)}_{n}=1\otimes\ldots\otimes p_{n}\otimes\ldots\otimes 1$ where $p_{n}$ is in the $i$th position. There is another interesting basis of $\Lambda_{t_{1},t_{2}}^{\otimes l}$ given by the wreath Macdonald polynomials of Haiman [Haiman1], see also the survey of Orr-Shimozono [OrrShim]. We will review the basics of wreath Macdonald theory in section 2. Under the isomorphism (1), the skyscraper sheaf of a fixed point $\lambda$ is mapped to the wreath Macdonald polynomial $H_{\lambda}$.
@@
This theorem closely resembles the wreath Cauchy identity. This is not a coincidence; rather, we will use the wreath Cauchy identity in the proof. We can equivalently recast the theorem as giving an interpretation to a special evaluation of wreath Macdonald polynomials.
@@
Setting $z_{i}=0$, we recover the evaluation formula for wreath Macdonald polynomials, see Proposition 2.7. When the core is empty, this states
$H_{\lambda}|_{p^{(j)}_{k}=\delta_{0,j}-\delta_{m,j}(-u)^{k}}=\prod_{\begin{subarray}{c}\square=(a,b)\in\lambda\\ a-b\equiv m\mod l\end{subarray}}(1+uq^{b-1}t^{a-1})$
This formula, and its generalization to arbitrary core, was recently proven by Romero and Wen in [WenRomero].
@@ 1.3 Ingredients in the proof
Let us conclude the introduction by discussing the ingredients in the proof of Theorem 1.1.
Viewing $\mathcal{M}(n)$ as a Nakajima quiver variety not only allows one to define quasimap moduli spaces but also opens the door to the representation-theoretic machinery developed in [Ok17] and [OkSm]. In particular, [Ok17] and [OkSm] construct an action of a Hopf algebra on the $K$-theory of quiver varieties and prove that the capping operator is determined by $q$-difference equations involving operators in this Hopf algebra. The important recent work of Tianqing Zhu [zhu2] identifies this Hopf algebra with the quantum toroidal algebra of type $A$, confirming a longstanding expectation. The cohomological version of this identification was proven earlier by Botta-Davison [BD] and Schiffman-Vasserot [SV23] independently. By Zhu’s theorem, we are able to use known results about quantum toroidal algebras and their slope subalgebras to approach capped vertex functions. Studying the $R$-matrix of the slope zero subalgebra in the tensor product of two Fock representations is what leads us to restrict our attention to $\tau_{0}$ rather than a general $\tau_{m}$.
The quiver variety perspective on $\mathcal{M}(n)$ is again useful, as one can vary the “framing” vector. Viewed as a Nakajima quiver variety, the dimension vector for $\mathcal{M}(n)$ is $\mathsf{v}=(n,n,\ldots,n)$ and the framing vector is $\mathsf{w}=(1,0,\ldots,0)=:\delta_{0}$. As is well-known, see [Naktens], one can embed any quiver variety as a $\mathbb{C}^{\times}$-fixed component of another quiver variety with larger framing. The gauge group $G_{\mathsf{v}}$, and hence the space of descendants $\mathcal{D}$, is left unchanged. Okounkov proves that for any given $\tau$, there exists a sufficiently large framing vector such that the quantum corrections of $\hat{V}^{\left(\tau\right)}$ vanish, see [Ok17] section 7. We refine Okounkov’s result here, proving the following.

### v3

# Wreath Macdonald polynomials, quiver varieties, and quasimap counts

@@ Abstract
We study the $K$-theoretic enumerative geometry of cyclic Nakajima quiver varieties, with particular focus on $\text{Hilb}^{n}([\mathbb{C}^{2}/\mathbb{Z}_{l}])$, the equivariant Hilbert scheme of points on $\mathbb{C}^{2}$. The direct sum over $n$ of the equivariant $K$-theories of these varieties is known to be isomorphic to the ring symmetric functions in $l$ colors, with structure sheaves of torus fixed points identified with wreath Macdonald polynomials. Using properties of wreath Macdonald polynomials and the recent identification of the Maulik-Okounkov quantum affine algebra for cyclic quivers with the quantum toroidal algebras of type $A$, we derive an explicit formula for the generating function of capped vertex functions of $\text{Hilb}^{n}([\mathbb{C}^{2}/\mathbb{Z}_{l}])$ with descendants given by exterior powers of the 0th tautological bundle. We also sharpen the large framing vanishing results of Okounkov, providing a class of descendants and cyclic quiver varieties for which the capped vertex functions are purely classical. Our results also suggest certain integrality and wall-crossing conjectures for capped vertex functions.
@@ 1.1 Enumerative geometry of quiver varieties
The pioneering work of Maulik-Okounkov [35] demonstrated that the quantum cohomology of Nakajima quiver varieties is controlled by certain geometrically constructed quantum groups. The $K$-theoretic version of this work, relating quasimaps counts to quantum affine algebras, was developed in [46] and [48]. This interplay between geometry and representation theory has lead to a large body of work in areas like symmetric functions, integrable systems, characteristic classes, and $q$-difference equations, [1, 2, 7, 24, 31, 47, 48, 52, 53, 58, 68]. On the one hand, $K$-theoretic counts present additional technical challenges compared to cohomological counts; on the other hand, they are required to observe additional structure, most importantly 3d mirror symmetry [1, 2, 11, 14, 18, 60]. By the degeneration and glue formulas [46], general $K$-theoretic quasimap counts in Nakajima quiver varieties are built from the following three objects: the (bare) vertex, the capping operator, and the capped vertex. Each corresponds to a different choice of “boundary condition” for quasimaps. We touch briefly on the the first two before turning in more detail to the third, which is the main focus of this paper.
@@
For many examples of interest, the bare vertex can be computed by localization, see [2, 18]. It provides a class of hypergeometric series generalizing the well-known $_{n+1}\phi_{n}$ basic hypergeometric series studied at length in, for example, [22]. Vertex functions are known to satisfy certain $q$-difference equations with regular singularities. The monodromy of these equations was determined in [1] and [47].
@@
The capping operator is the fundamental solution of the quantum difference equation, which is the $K$-theoretic version of the quantum differential equation. For Nakajima quiver varieties, these equations generalize the qKZ and dynamical difference equations studied in [20, 61, 62]. The geometric approach to these equations was used by Aganagic-Okounkov in [2] to provide integral formulas for solutions.
@@
The third fundamental curve count, the capped vertex, is the main object of interest for this paper. In some ways, the capped vertex is simpler than the first two curve counts: it is defined in non-localized $K$-theory and is purely classical for quiver varieties with large enough framing, a phenomenon known as “large framing vanishing” and proved in [46]. It was proved by Smirnov [59] for the Jordan quiver and by Zhu [69] for general quivers (and earlier in cohomology in [50]) that the capped vertex is a rational function of the quantum parameters (typically called Kähler parameters in this setting). On the other hand, computing the capped vertex explicitly in examples is difficult; the localization theorem essentially reduces the problem to computing both the bare vertex and capping operator. Building on [59], the recent work of Smirnov and the first author [3] developed explicit formulas for the capped vertex for what is arguably the most important quiver variety: the Hilbert scheme of points in the plane. Their formulas reveal a beautiful relationship with Macdonald polynomials and give an enumerative-geometric interpretation to certain evaluations of them.
@@
In this paper, we push the methods of Ayers-Smirnov further, studying the capped vertex functions for cyclic quiver varieties. Our main theorem, Theorem 1.1 below, describes the capped vertex functions for certain cyclic quiver varieties in terms of wreath Macdonald polynomials [5, 25]. Our main theorem can be interpreted as describing capped vertex functions as certain evaluations of wreath Macdonald polynomials, see Theorem 1.2.
@@ 1.2 The capped descendant vertex
Let $\mathcal{M}:=\text{Hilb}^{n}([\mathbb{C}^{2}/\Gamma_{l}])$ where $\Gamma_{l}=\mathbb{Z}/l\mathbb{Z}$, which acts on $\mathbb{C}^{2}$ by $\zeta\cdot(x,y)=(\zeta x,\zeta^{-1}y)$ where $\zeta$ is a primitive $l$th root of unity.. Concretely, the variety $\mathcal{M}$ parameterizes $\Gamma_{l}$ invariant ideals $I\subset\mathbb{C}[x,y]$ such that $\text{char}_{\Gamma_{l}}\left(\mathbb{C}[x,y]/I\right)=\sum_{i=0}^{l-1}n\chi^{i}$, where $\chi^{i}$ are the irreducible characters of $\Gamma_{l}$. In particular, $\mathcal{M}\subset\text{Hilb}^{nl}(\mathbb{C}^{2})$.
The variety $\mathcal{M}$ is an example of a Nakajima quiver variety [40, 41] with dimension vector $\mathsf{v}=(n,n,\ldots,n)$ and thus has a canonical presentation as a geometric invariant theory quotient
@@
This description is reviewed in section 3. Hence we can apply the general machinery of quasimaps to GIT quotients developed in [9]. To that end, let $\text{QM}^{d}$ be the moduli space of stable quasimaps of degree $d\in\mathbb{Z}^{l}$ relative to the point $0\in\mathbb{P}^{1}$. It compactifies the space of genuine maps from $\mathbb{P}^{1}$ to $\mathcal{M}$.
@@
The notation $\mu^{-1}(0)/G_{\mathsf{v}}]$ in the bottom right corner of the diagram refers to the stack quotient. By definition of $\mathcal{M}$, there is an inclusion $\iota:\mathcal{M}\hookrightarrow\mathcal{M}^{\text{st}}$. There is an action of a torus $\mathsf{T}$ on each piece of the above diagram such that the evaluation maps are $\mathsf{T}$-equivariant.
@@
The series $\hat{V}^{\left(\tau\right)}$ was proved in [59] to be a rational function in $z$ for the rank $r$ instanton moduli space, for which the cyclic quiver at $l=1$ is the rank 1 version. An explicit formula for the $l=1$ case was given in [3]. Quasimap counts in cyclic quiver varieties are related to curve counts in threefolds, see [32]. The cohomological version of this relationship goes back to Okounkov-Pandharipande [45] and Maulik-Oblomkov [34].
@@
where the superscript Sym refers to Laurent polynomials that symmetric in the Chern roots $x_{i,1},\ldots,x_{i,n}$ for each $i$ separately. Equivalently, $\mathcal{D}$ is the ring generated by the Schur functors of the tautological vector bundles $\mathcal{V}_{i}$, $0\leq i\leq l-1$, on $\mathcal{M}$. Let $\tau_{i}:=\sum_{j=0}^{n}\bigwedge^{j}\mathcal{V}_{i}$.
@@
Now we consider the varieties $\mathcal{M}$ for all $n$ and denote $\mathcal{M}(n):=\text{Hilb}^{n}([\mathbb{C}^{2}/\Gamma_{l}])$. By [5], there is an isomorphism
@@
where $\Lambda_{t_{1},t_{2}}=\mathbb{Q}(t_{1},t_{2})[p_{1},p_{2},\ldots]$ is the ring of symmetric functions [^2] The usual $q$ and $t$ of Macdonald theory match our parameters by $q=t_{1},t=t_{2}$.. The ring $\Lambda_{t_{1},t_{2}}^{\otimes l}$ is generated as an algebra over $\mathbb{Q}(t_{1},t_{2})$ by the “colored” power sums: $p^{(i)}_{n}=1\otimes\ldots\otimes p_{n}\otimes\ldots\otimes 1$ where $p_{n}$ is in the $i$th position. There is another interesting basis of $\Lambda_{t_{1},t_{2}}^{\otimes l}$ given by the wreath Macdonald polynomials of Haiman [25], see also the survey of Orr-Shimozono [49]. We will review the basics of wreath Macdonald theory in section 2. Under the isomorphism (1), the skyscraper sheaf of a fixed point $\lambda$ is mapped to the wreath Macdonald polynomial $H_{\lambda}$. The two spaces in (1) are equipped with the action of the quantum toroidal algebra, compatible up to certain scalars. We assume in this paper, see section 6.1, that these scalars can be neglected.
@@
Theorem 1.1 closely resembles the Cauchy identity for wreath Macdonald polynomials. This is not a coincidence; rather, we will use the wreath Cauchy identity in the proof. We can equivalently recast the theorem as giving an interpretation to a special evaluation of wreath Macdonald polynomials.
@@
Setting $z_{i}=0$ in Theorem 1.2, we recover the evaluation formula for wreath Macdonald polynomials, see Proposition 2.7. When the core is empty, this states
This formula, and its generalization to arbitrary core, was recently proved by Romero and Wen in [54].
@@
We deduce a few corollaries from Theorem 1.1.
Corollary 1.3.
Viewed as a rational function of $z$, the poles of $\hat{V}^{\left(\bigwedge^{j}\mathcal{V}_{0}\right)}$ are contained in the locus $\{(-z_{0}\ldots z_{l-1}\hbar^{-1/2})^{k}=1\,\mid\,1\leq k\leq j\}$.
Specializing $z_{0}\ldots z_{l-1}=-\hbar^{-1/2}$ yields the following.
Corollary 1.4.
We have the evaluation
$\hat{V}^{\left(\tau_{0}\right)}|_{z_{0}\ldots z_{l-1}=-\hbar^{-1/2}}=1$
Equivalently,
$\hat{V}^{\left(\bigwedge^{j}\mathcal{V}_{0}\right)}|_{z_{0}\ldots z_{l-1}=-\hbar^{-1/2}}=0\quad\text{for $j>0$}$
Consider the following automorphisms on the ring of symmetric functions
$\mathsf{swap}$ $:t_{1}\mapsto t_{2},\,t_{2}\mapsto t_{1}$
$\mathsf{neg}$ $:p^{(i)}_{n}\mapsto p^{(-i)}_{n}$
where the superscripts are taken modulo $l$. It is known that $\mathsf{swap}\,\mathsf{neg}\,H_{\lambda}=H_{\lambda^{t}}$ when $\text{core}_{l}(\lambda)=\emptyset$, see [49] Proposition 3.21. One readily checks that $\mathsf{swap}\,C^{(i)}_{n}=C^{(-i)}_{n}$ and $\mathsf{swap}\,N_{\lambda}=N_{\lambda^{t}}$. So applying $\mathsf{swap}\,\mathsf{neg}$ to Theorem 1.1 gives the following.
Corollary 1.5.
$\mathsf{swap}\,\hat{V}^{\left(\tau_{0}\right)}_{\lambda}=\hat{V}^{\left(\tau_{0}\right)}_{\lambda^{t}}$
@@ 1.3 Ingredients in the proof
Let us now discuss the ingredients in the proof of Theorem 1.1.
Viewing $\mathcal{M}(n)$ as a Nakajima quiver variety not only allows one to define quasimap moduli spaces but also opens the door to the representation-theoretic machinery developed in [46] and [48]. In particular, [46] and [48] construct an action of a Hopf algebra on the $K$-theory of quiver varieties and prove that the capping operator is determined by $q$-difference equations involving operators in this Hopf algebra. The important recent work of Zhu [68] identifies this Hopf algebra with the quantum toroidal algebra of type $A$, confirming a longstanding expectation. The cohomological version of this identification was proved earlier by Botta-Davison [6] and Schiffman-Vasserot [56] independently. By Zhu’s theorem, we are able to use known results about quantum toroidal algebras and their slope subalgebras to approach capped vertex functions. Studying the $R$-matrix of the slope zero subalgebra in the tensor product of two Fock representations is what leads us to restrict our attention to $\tau_{0}$ rather than a general $\tau_{m}$.
The quiver variety perspective on $\mathcal{M}(n)$ is again useful, as one can vary the “framing” vector. Viewed as a Nakajima quiver variety, the dimension vector for $\mathcal{M}(n)$ is $\mathsf{v}=(n,n,\ldots,n)$ and the framing vector is $\mathsf{w}=(1,0,\ldots,0)=:\delta_{0}$. As is well-known, see [42], one can embed any quiver variety as a $\mathbb{C}^{\times}$-fixed component of another quiver variety with larger framing. The gauge group $G_{\mathsf{v}}$, and hence the space of descendants $\mathcal{D}$, is left unchanged. Okounkov proves that for any given $\tau$, there exists a sufficiently large framing vector such that the quantum corrections of $\hat{V}^{\left(\tau\right)}$ vanish, see [46] section 7. We refine Okounkov’s result here, proving the following.